OpenStudy (anonymous):

what is the equation of the circle with center (2, –5) that passes through the point (–2, 10)?

OpenStudy (anonymous):

general equation of a circle: \[(x - h)^{2} + (y - k)^{2} = r^2\] where the circle's center is the point (h, k) and the radius has length r

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

if the circle passes through (-2, 10), it means that the radius is the distance between the center and that point. \[d = r = \sqrt{(x_2 - x_1)^{2} + (y_2 - y_1)^{2}}\]

OpenStudy (anonymous):

d meaning distance for the standard distance formula above (not to be confused with diameter)

OpenStudy (anonymous):

So what should I do, I don't understand. I follow what your writing but I don't know how to solve.

OpenStudy (anonymous):

you simply plug in. lets call the center (x1,y1) and the point (x2,y2)

OpenStudy (anonymous):

so (x1,y1) is (2,-5) and then (x2,y2) is (-2,10). Then what do I do?

OpenStudy (anonymous):

yes, you'd find r: r = \[r = \sqrt{(-2 - 2)^{2} + (10 - (-5))^{2}} = \sqrt{16 + 225}\] then you have h, k and r for the general equation formula. where (h,k) = (x1,y1) = center of circle \[(x - 2)^{2} + (y - (-5))^{2} = (\sqrt{16+225})^{2}\] simplifying will give the final answer; the equation of the circle

OpenStudy (anonymous):

I am stuck...

OpenStudy (anonymous):

I don't know how to simplify that..

OpenStudy (anonymous):

simplified version would be: \[(x - 2)^{2} + (y + 5)^{2} = 241\]

OpenStudy (anonymous):

I am in all seriousness don't know how. I am not trying to just get the answer. I want to know how to do it.

OpenStudy (anonymous):

i believe :). i didn't suspect you of trying to just get the answer lol

OpenStudy (anonymous):

How is that the answer it's not a choice do I have to do something else

OpenStudy (anonymous):

I was just wanting to make sure you knew.

OpenStudy (whpalmer4):

what are your answer choices?

OpenStudy (whpalmer4):

Are \(x^2+y^2-4x+10y+29=241\) or \(x^2+y^2-4x+10y=212\) among your answer choices? That's what you would get if you expanded the answer.

OpenStudy (whpalmer4):

Though normally one would keep it in the form given by Euler271, as it is easy to then see the properties of the circle without laboriously graphing it.